# The Uli Dataset: An Exercise in Experience Led Annotation of oGBV

### v2

@@ 1 Introduction
This dataset, inspired from values of feminist technologies such as inclusion, intersectionality and care, is an attempt at participatory models of machine learning development Clancy (2021). The definition of gendered abuse, as well as the annotations, came from activists and researchers who identify as a marginalized gender and have encountered or responded to online or offline abuse. This paper describes the process of creation of the dataset in three Indic languages: Hindi, Tamil and Indian English.
@@ 3 Related Work
While most of the aforementioned papers focus on English language content, there has been a push to expand abuse and hate speech detection to languages other than English. Within Indic languages, Hindi has received considerable attention. Mandl et al. (2019, 2020b) proposed a dataset for hate speech in Hindi language consisting of $5K$ and $6K$ posts sourced from Twitter. Following the previous work, Mandl et al. (2020a) shared another hate speech dataset of $3.6K$ posts scrapped from YouTube and Twitter. Bohra et al. (2018) introduced a code-mixed Hindi dataset on Hate Speech containing $4.5K$ tweets, out of which, $1.6K$ tweets are labelled hateful, and the remaining $2.9K$ are non-hateful. Tweets are annotated as hate speech or normal speech. Saroj and Pal (2020) proposed a dataset for Hindi language on offensive speech containing $2K$ posts from Twitter and Facebook. Velankar et al. (2021) created a dataset for Hindi and Marathi on hate and offensive speech with $4.5K$ and $2K$ posts respectively. Romim et al. (2021) created a dataset on hate speech consisting of $30K$ comments in Bengali from YouTube and Facebook, $10K$ comments are annotated as hateful. Gupta et al. (2022) proposed a large-scale (150K) abusive speech dataset of comments in Hindi, Tamil, Telugu, Kannada and Malayalam sourced from ShareChat.  Chakravarthi et al. (2021) created a code-mixed dataset on offensive speech consisting of YouTube comments in Kannada, Malayalam, and Tamil, with 7.7K comments for Kannada, 20K for Malayalam, and 43K for Tamil. Bhardwaj et al. (2020) collect posts from Twitter and Facebook in Hindi and provide annotations for hostile posts including fake news, hate speech, and other offensive posts. The  Kumar et al. (2021) dataset specifically looked at gendered and communally charged comments in four Indian languages. The dataset was annotated at three levels: aggression, gender bias, and communal bias.
@@ 4.2 Stratified Pooling
Our annotation budget determined the dataset size: roughly 8000 posts in three languages. Despite our strategy of collecting data based on problematic keywords, the majority of the dataset was non-oGBV. Creating an annotation set by randomly sampling posts from the larger unlabelled dataset of 1.3 million tweets would have resulted in a very small dataset of tweets containing gendered abuse. Thus, we used stratified pooling to create a dataset in which the percentage of abuse is higher than the larger data corpus, and possibly higher than Twitter in general. To do this, we first assign noisy labels to our unlabelled dataset using democratic co-training as done in prior work (Rosenthal et al., 2021). We used various models trained on open-source datasets for related tasks of offensive language detection, misogyny detection, and hate speech detection. In Table 1, we show the set of datasets and models used for our semi-supervised annotation per language.
@@ 10 Model performance
To assess performance of existing approaches to detect oGBV, we tested models from the automated abuse detection literature on our dataset. Specifically, we created train and test sets from annotations for Label 2 and fine-tuned models on them. We considered all data annotated by a single annotator as training data and ones annotated by multiple people (Table  5) as test data, using majority labelling for the final label. For the models, we used IndicBERT (Kakwani et al., 2020), which is a multilingual ALBERT model trained on Indic language data, and XLM-T (Barbieri et al., 2022), a RoBERTa model fine-tuned on Twitter data, on our dataset. Our choice of model was motivated by the strong performance of IndicBERT over other multilingual models like mBERT and XLM-R when evaluated on tasks in Indian languages (Kakwani et al., 2020). For XLM-T, we relied on its pre-training on Twitter data and strong performance in prior social media based datasets as our primary motivation for inclusion. For all the models, we trained for 5 epochs with a learning rate of 5e-06, a batch size of 8 and the Adam optimizer. To avoid overfitting, we implement an early stopping mechanism conditioned on the evaluation F1 macro with a patience of 5 steps. We report the result in Table 8.

### v3

@@ 1 Introduction
This dataset, inspired from values of feminist technologies such as inclusion, intersectionality and care, is an attempt at participatory models of machine learning development Clancy (2021). The definition of gendered abuse, as well as the annotations, came from activists and researchers who identify as a marginalized gender and have encountered or responded to online or offline abuse. This paper describes the process of creation of the dataset in three Indian languages: Hindi, Tamil and Indian English.
@@ 3 Related Work
While most of the aforementioned papers focus on English language content, there has been a push to expand abuse and hate speech detection to languages other than English. Within Indian languages, Hindi has received considerable attention. Mandl et al. (2019, 2020b) proposed a dataset for hate speech in Hindi language consisting of $5K$ and $6K$ posts sourced from Twitter. Following the previous work, Mandl et al. (2020a) shared another hate speech dataset of $3.6K$ posts scrapped from YouTube and Twitter. Bohra et al. (2018) introduced a code-mixed Hindi dataset on Hate Speech containing $4.5K$ tweets, out of which, $1.6K$ tweets are labelled hateful, and the remaining $2.9K$ are non-hateful. Tweets are annotated as hate speech or normal speech. Saroj and Pal (2020) proposed a dataset for Hindi language on offensive speech containing $2K$ posts from Twitter and Facebook. Velankar et al. (2021) created a dataset for Hindi and Marathi on hate and offensive speech with $4.5K$ and $2K$ posts respectively. Romim et al. (2021) created a dataset on hate speech consisting of $30K$ comments in Bengali from YouTube and Facebook, $10K$ comments are annotated as hateful. Gupta et al. (2022) proposed a large-scale (150K) abusive speech dataset of comments in Hindi, Tamil, Telugu, Kannada and Malayalam sourced from ShareChat.  Chakravarthi et al. (2021) created a code-mixed dataset on offensive speech consisting of YouTube comments in Kannada, Malayalam, and Tamil, with 7.7K comments for Kannada, 20K for Malayalam, and 43K for Tamil. Bhardwaj et al. (2020) collect posts from Twitter and Facebook in Hindi and provide annotations for hostile posts including fake news, hate speech, and other offensive posts. The  Kumar et al. (2021) dataset specifically looked at gendered and communally charged comments in four Indian languages. The dataset was annotated at three levels: aggression, gender bias, and communal bias.
@@ 4.2 Stratified Pooling
Our annotation budget determined the dataset size: roughly 8000 posts in three languages. Despite our strategy of collecting data based on problematic keywords, the majority of the dataset was non-oGBV. Creating an annotation set by randomly sampling posts from the larger unlabelled dataset of 1.3 million tweets would have resulted in a very small dataset of tweets containing gendered abuse. Thus, we used stratified pooling to create a dataset in which the percentage of abuse is higher than the larger data corpus, and possibly higher than Twitter in general. To do this, we first assign noisy labels to our unlabelled dataset using democratic co-learning as done in prior work (Rosenthal et al., 2021). We used various models trained on open-source datasets for related tasks of offensive language detection, misogyny detection, and hate speech detection. In Table 1, we show the set of datasets and models used for our semi-supervised annotation per language.
@@ 10 Model performance
To assess performance of existing approaches to detect oGBV, we tested models from the automated abuse detection literature on our dataset. Specifically, we created train and test sets from annotations for Label 2 and fine-tuned models on them. We considered all data annotated by a single annotator as training data and ones annotated by multiple people (Table  5) as test data, using majority labelling for the final label. For the models, we used IndicBERT (Kakwani et al., 2020), which is a multilingual ALBERT model trained on Indian language data, and XLM-T (Barbieri et al., 2022), a RoBERTa model fine-tuned on Twitter data, on our dataset. Our choice of model was motivated by the strong performance of IndicBERT over other multilingual models like mBERT and XLM-R when evaluated on tasks in Indian languages (Kakwani et al., 2020). For XLM-T, we relied on its pre-training on Twitter data and strong performance in prior social media based datasets as our primary motivation for inclusion. For all the models, we trained for 5 epochs with a learning rate of 5e-06, a batch size of 8 and the Adam optimizer. To avoid overfitting, we implement an early stopping mechanism conditioned on the evaluation F1 macro with a patience of 5 steps. We report the result in Table 8.
@@ 12 Limitations
13 Acknowledgments
We would like to thank the funders who supported the project: Digital Society Challenge award granted to the Center for Internet and Society and Tattle Civic Tech; Mozilla Foundation’s Data Futures Lab Cohort; and in-kind support by GitHub. At different stages of the project we consulted with experts in the field of gender studies, machine learning, content moderation and computational social science. The number of experts are too many to list here but we are grateful for their time and advice.